OpenStudy (anonymous):

two snails are 374 centimeters apart. They crawl toward each other at rates that differ by 5 cm/min. after 22 minutes they meet. how fast is each snail crawling?

OpenStudy (anonymous):

let

OpenStudy (kropot72):

Let speed of slower snail = s Then speed of faster snail = s + 5 Distance travelled by slower snail = s * 22 cm Distance travelled by faster snail = (s + 5) * 22 cm Combined distance travelled by snails = 374 = 22s + 22s + 110 Now you can solve to find s.